Grade 12th passPhysical Chemistry

Sample 1 has higher o^18/o^16 ratio of atoms are compared to sample 2 . The average mass of oxygen insample 1 is............ That of sample 2.

Harpreet Singh
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Total particles in 1 mole of oxygen = 6.023 moles.' 
Nuclear number of oxygen is 8 , atomic weight is 16 . 
Test 1 has oxygen as 18/16 = 9/8 proportion . 
In this manner normal mass of both the proportion is = 9/8 x 16 = 18 
In this manner test 1 has 2 more oxygen iotas than the main specimen.